BAYESIAN CALIBRATION IN THE ESTIMATION OF THE AGE OF RHINOCEROS

J. L. DU PLESSIS AND A. J. VAN DER MERWE

Department of Mathematical Statistics, Faculty of Science, University of the O.F.S., P.O. Box 339, Bloemfontein, 9300, Republic of South Africa

(Received June 28, 1993; revised April 11, 1995)

Abstract. In this paper the Bayesian approach for nonlinear multivariate calibration will be illustrated. This goal will be achieved by applying the Gibbs sampler to the rhinoceros data given by Clarke (1992, Biometrics, 48(4), 1081– 1094). It will be shown that the point estimates obtained from the profile likelihoods and those calculated from the marginal posterior densities using improper priors will in most cases be similar.

Key words and phrases: Anterior horn length, multivariate Bayesian calibration, nonlinear response model, credibility intervals, Gibbs sampler, posterior horn length, posterior distribution, rhinoceros data, Student-t family.